OpenStudy (anonymous):

In isosceles trapezoid JKLM, Leg JK = 5x + 10, base KL = 6x + 2, and leg LM = 2x +8. Find the value of x. 1) -2/3 2) 3/2 3) 6 4) -12

OpenStudy (mertsj):

The legs are equal. 5x+10=2x+8 3x=-2 x=-2/3

OpenStudy (anonymous):

Thank you soo much